OpenStudy (anonymous):

w^4-20w^2-2=0 w^4-20w^2-2=0 @Mathematics

OpenStudy (agreene):

\[w=\pm\sqrt{10+\sqrt{102}}\] and \[w=\pm (i\sqrt{\sqrt{102}-10})\]

jhonyy9 (jhonyy9):

w squared note x and will get x2 -20x -2=0 x_1,_2=(20+/- sqrt(400+8))/2 =(20+/- sqrt408)/2 =10+/- sqrt102 than w2=10+/- sqrt102 so w_1= - sqrt(10+/-sqrt102) and w_2 = sqrt(10+/-sqrt102)

jhonyy9 (jhonyy9):

yes right w_1 sign w_1,_3 and w_2 sign w_2,_4

jhonyy9 (jhonyy9):

and hence there are those 4 solutions